OpenStudy (anonymous):

find the distance from the line y-6=m(x+1) to the point given below

OpenStudy (anonymous):

\[(\sqrt{(1+m^2)},6+\sqrt{(1+m^2)} )\]\[(1,4)\]and \[(-2,1)\]these are the three given points

OpenStudy (amistre64):

what methods do you have available at your disposal?

OpenStudy (anonymous):

I seeing this formula but not sure what to do..\[d=\frac{ Ax+By+C }{ \pm \sqrt{A^2+B^2} }\]

OpenStudy (amistre64):

i cant say i know what that refers to either ... but i have an idea on how to find a solution. we need to know the distance formula, how to determine a perpendicular slope, formulate a line equation from a point and a slope, and determine where 2 lines cross at.

OpenStudy (anonymous):

ok how do we do that

OpenStudy (amistre64):

another idea involves making a circle at a point, seeing where the line crosses it, and taking the midpoint, then distancing the results

OpenStudy (amistre64):

well, given a slope of a line: m -1/m creates a perpendicular slope ... for a given point (a,b), the form of the perp line is: -1/m(x-a) +b finding where the two lines meet ... is equating them -1/m (x-a) + b = m(x+1)+6 -x/m +a/m + b = mx +m +6 a/m + b -(m +6) = mx+x/m a/m + b -(m +6) = x(m+1/m) (a/m + b -(m +6))/(m+1/m) = x knowing x, we find y by substitution etc ...

OpenStudy (amistre64):

but if you can figure out that formula, and it is one for your course ... go for it :)

OpenStudy (amistre64):

 i think that formula of yours is from the distance between 2 planes tho

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

we don't know whats m

OpenStudy (amistre64):

m is just a general value ... the solution will be general as well

OpenStudy (anonymous):

so y=(a/m + b -(m +6))/(m+1/m) ??

OpenStudy (amistre64):

given y = m(x+1)+6 and x = (a/m + b -(m +6))/(m+1/m) , i hope i did that right :) y = mx + m + 6 mx = (a + bm -m^2 +6m)/(m+1/m) = -m(m^2- (6+b)m -a)/(m^2+1) y = -m(m^2- (6+b)m -a)/(m^2+1) + m + 6 might be prudent to find the values for x and y before trying to process a general formula, it is starting to look messy

OpenStudy (amistre64):

lets use the (1,4) for (a,b)

OpenStudy (amistre64):

and for simplicity, assume a slope of 1 for brevity

OpenStudy (anonymous):

ok so we can use any poits for (a,b)?

OpenStudy (amistre64):

x = (a/m + b -(m +6))/(m+1/m) x = (1/1 + 4 -(1 +6))/(1+1/1) x = (-2)/(2) = -1 y = m(-1+1)+6 = 6 well, (a,b) should be a stated point, since that is the one we are trying to define a line with and cross it

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D1%28x%2B1%29%2B6%2C+y%3D-1%28x-1%29%2B4 the good news maybe that our formulas for finding the crossing point are good

OpenStudy (anonymous):

I think ill understand this ...if I wasn't this sleepy

OpenStudy (amistre64):

what is the distance from (1,4) to (-1,6) ?

OpenStudy (anonymous):

4+4=sq.rt 8

OpenStudy (amistre64):

so 2 sqrt2

OpenStudy (amistre64):

but this assumes a slope; m=1

OpenStudy (amistre64):

trying to think of a vector way to approach it, or translation of the line so that we are finding the distance from the origin

OpenStudy (anonymous):

ok if it work....but this is geometry though

OpenStudy (amistre64):

we can use trig?

OpenStudy (anonymous):

we can use anything ..I guess

OpenStudy (amistre64):

well, y-6 = m(x+1) can be dropped 6 units to pass thru the origin as y = m(x+1), since we drop the line, we drop the point by 6 as well (1,4-6) = (1,-2)